# On the Performance Analysis of Momentum Method: A Frequency Domain Perspective

### v1

@@
On the Performance Analysis of Momentum Method: A Frequency Domain Perspective
Xianliang Li1,2   Jun Luo1,211footnotemark: 1   Zhiwei Zheng3   Hanxiao Wang2,4
Li Luo5   Lingkun Wen2,6   Linlong Wu7   Sheng Xu1
1Shenzhen Institute of Advanced Technology  2University of Chinese Academy of Sciences
3University of California, Berkeley  4Institute of Automation  5Sun Yat-sen University
6Shanghai Astronomical Observatory  7University of Luxembourg
yinleung.ley@gmail.com, j.luo3@siat.ac.cn, zhiwei.zheng@berkeley.edu
wanghanxiao18@mails.ucas.ac.cn, luoli33@mail2.sysu.edu.cn, wenlingkun@shao.ac.cn
linlong.wu@uni.lu, sheng.xu@siat.ac.cn
Equal contribution.Corresponding author.
Momentum-based optimizers are widely adopted for training neural networks. However, the optimal selection of momentum coefficients remains elusive. This uncertainty impedes a clear understanding of the role of momentum in stochastic gradient methods. In this paper, we present a frequency domain analysis framework that interprets the momentum method as a time-variant filter for gradients, where adjustments to momentum coefficients modify the filter characteristics. Our experiments support this perspective and provide a deeper understanding of the mechanism involved. Moreover, our analysis reveals the following significant findings: high-frequency gradient components are undesired in the late stages of training; preserving the original gradient in the early stages, and gradually amplifying low-frequency gradient components during training both enhance generalization performance. Based on these insights, we propose Frequency Stochastic Gradient Descent with Momentum (FSGDM), a heuristic optimizer that dynamically adjusts the momentum filtering characteristic with an empirically effective dynamic magnitude response. Experimental results demonstrate the superiority of FSGDM over conventional momentum optimizers.
@@ 1 Introduction
Momentum has achieved great success in deep learning applications when combined with Stochastic Gradient Descent (SGD) (Robbins and Monro,, 1951). Among various momentum methods (Polyak,, 1964; Nesterov,, 1983; Van Scoy et al.,, 2017; Ma and Yarats,, 2018; Kidambi et al.,, 2018), one of the most prevalent variants is the momentum method utilized within Stochastic Gradient Descent with Momentum (SGDM) (Sutskever et al.,, 2013; Paszke et al.,, 2019), which can be expressed as:
@@
where $g_{t}$ denotes the gradient at iteration $t$, ${m}_{t}$ is the momentum buffer, and $x_{t}$ represents the learnable parameters. The momentum coefficients $u_{t}$ and $v_{t}$ control the influence of the previous momentum and the current gradient, respectively, and $\alpha_{t}$ is the learning rate. For these time-variant momentum coefficients, a multistage setting has been commonly adopted in the machine learning community (Aybat et al.,, 2019; Kulunchakov and Mairal,, 2019; Liu et al.,, 2020). Throughout this paper, we refer to this formulation, which decouples the two momentum coefficients, as Standard-SGDM. In contrast, another prevalent variant couples the two momentum coefficients using the Exponential Moving Average (EMA) method (Gardner Jr,, 1985), leading to the formulation of EMA-SGDM:
@@
where $u_{t}\in[0,1)$ is the momentum coefficient. Notably, this coupled momentum formulation is a special case of the decoupled one, i.e., Standard-SGDM with $v_{t}=1-u_{t}$. Our experiments show performance gaps between these two formulations. Moreover, how the momentum coefficients change over time can significantly affect the test accuracy (see Section 3). The existence of these two distinct momentum formulations and their differing performances raises two primary questions in modern deep learning:
@@
For Question 1, some literatures have investigated the convergence of the coupled method (Mai and Johansson,, 2020; Li et al.,, 2022). Liu et al., (2020) argued that coupling the coefficients leads only to a constant scaling difference. Wang et al., (2024) further demonstrated that the mathematical equivalence between EMA-SGDM and Standard-SGDM can be achieved by adjusting the momentum coefficients and the learning rates in a coupled way. However, in practice, learning rate schedules are typically independent of momentum coefficient tuning during network training. On the other hand, popular frameworks like PyTorch (Paszke et al.,, 2019) adopt a decoupled momentum strategy by default. In our framework, we tackle the first question from the frequency domain perspective, revealing the relationship between the coupled and decoupled constructions.
@@
Regarding Question 2, prior research offered diverse opinions on how the momentum coefficients should vary over time. Some studies preferred fixed decoupled momentum coefficients (Yan et al.,, 2018; Liu et al.,, 2018; Yu et al.,, 2019), commonly selecting $u_{t}$ values as 0.9 and $v_{t}$ value as 1. Liu et al., (2020) highlighted the benefits of stagewise learning rate schedules in EMA-SGDM, noting that $u_{t}$ can either remain constant or increase along with the stagewise adjustments. Conversely, Smith, (2018) demonstrated that decreasing the momentum coefficients while increasing the learning rate improves test performance. Moreover, Adaptive momentum methods (Kingma and Ba,, 2014; Reddi et al.,, 2018; Luo et al.,, 2019; Chen et al.,, 2018) proved the convergence of decreasing coupled momentum coefficients in the context of online convex optimization. Nonetheless, a consensus regarding the optimal time-variant pattern of the momentum coefficients has yet to be reached.
@@
To answer these questions, one has to understand how the momentum method affects the training process. Goh, (2017) analyzed the momentum method from the aspect of convergence and dynamics. Several prior studies (Cutkosky and Orabona,, 2019; Ma and Yarats,, 2018) speculated that averaging past stochastic gradients through momentum might reduce the variance of the noise in the parameter update, thus making the loss decrease faster. Polyak, (1964); Rumelhart et al., (1986) argued that the EMA momentum can cancel out oscillations along high-curvature directions and add up contributions along low-curvature directions. From the signal processing perspective, the EMA method acts as a discrete low-pass filter for smoothing out high-frequency fluctuations while retaining the low-frequency baseband pattern of the signal (Gardner Jr,, 1985). These points of view bring us a new insight into connecting the momentum update processes with the specific filters. In this aspect, the momentum methods with different coefficient selections can be interpreted in a unified frequency domain analysis framework, whereby Questions 1 and 2 are resolved.
@@
In this paper, we propose a novel frequency domain analysis framework to address the two questions and provide a deeper understanding of the role of momentum in stochastic optimization. To the best of our knowledge, this paper, for the first time, reveals the fundamental difference between Standard-SGDM and EMA-SGDM and uncovers the effects of the dynamic momentum coefficients clearly with the help of frequency domain analysis. Within our framework, high-frequency gradient components correspond to rapid fluctuations in gradient signals, while low-frequency gradient components represent smoother, gradual changes. Moreover, adjusting the momentum coefficients during training is equivalent to modifying the time-variant filtering characteristics of the momentum methods. This perspective not only explains the difference between various momentum methods but also provides practical guidelines for designing well-performed optimizers. Building upon these insights, we introduce the Frequency Stochastic Gradient Descent with Momentum (FSGDM) optimizer, which dynamically adjusts the momentum filter characteristics throughout training. Our experiments validate the effectiveness of FSGDM, showing that it outperforms conventional SGD-based momentum optimizers. The code will be released upon publication.
@@ 2 Frequency Domain Analysis Framework
This section introduces the background of Z-transform (Zadeh,, 1950) in signal processing and then proposes a new frequency domain analysis framework for momentum methods.
@@ 2.1 Z-Transform and Quasi-Stationary Approximation
Frequency analysis is a crucial technique for understanding how systems react to varying frequency components of input signals. Specifically, for discrete-time linear time-invariant systems, Z-transform is leveraged to examine how systems attenuate or amplify signals at specific frequencies, especially in the study of system stability, pole-zero behavior, etc. (Oppenheim et al.,, 1996).
@@
However, one key issue is that our momentum method can be inherently time-variant, as its coefficients may change stagewise throughout the training process. This variability makes it difficult to apply traditional Z-transform analysis. To overcome this, inspired by the Zadeh, (1961); Jury, (1964), we approximate the system as time-invariant in each discrete interval stage. By holding the momentum coefficients constant over every interval, we construct a time-invariant quasi-stationary system (Hubner and Tran-Gia,, 1991), enabling us to apply the Z-transform validly.
@@
where $\mu$ and $\nu$ are the increasing and decreasing factors  [^1] Note that different from the increasing sequence, the numerator of the decreasing sequence is $t+1$. This design avoids the zero gradients at the first training stage.. In Appendix B.1, we also examined the test set performance using other kinds of dynamic sequences. Under the above settings, for a given stage $k$ ($k=1,\cdots,N$), with $t\in[(k-1)\delta,k\delta-1]$, the momentum system  [^2] In this paper, momentum system refers to the momentum update process. becomes:
@@ 2.2 Frequency Domain Analysis of the Momentum Method
In this subsection, we introduce our frequency domain analysis framework and analyze the impacts of the momentum method on neural network training. We first apply Z-transform, denoted by $\mathcal{Z}$, to Equation (5):
@@
where $G(z)=\mathcal{Z}\{g_{t}\}$, $M(z)=\mathcal{Z}\{m_{t}\}$, and $z^{-1}M(z)=\mathcal{Z}\{m_{t-1}\}$. To obtain the frequency response of the momentum system during stage $k$, we evaluate the transfer function $H_{k}(z)$ on the unit circle (Oppenheim et al.,, 1996):
@@
Orthodox Momentum Systems are the ones whose amplitude of the magnitude response will not surpass $1$, like the EMA-SGDM (2). This kind of momentum system only shows attenuating characteristics. Specifically, the momentum system behaves as a low-pass filter when $u_{k}>0$ and a high-pass filter when $u_{k}<0$. Additionally, when $u_{k}$ gets close to 1, the momentum system will prefer to attenuate the gradient components with high frequencies. The visualization of the (dynamic) magnitude responses of orthodox momentum systems is in Section 3.1 and Appendix B.2.
For Unorthodox Momentum Systems where the amplitude of magnitude response will surpass $1$, such as selecting $u_{t}=0.9$ and $v_{t}=1$ in Standard-SGDM (1), the momentum system possesses both amplifying and attenuating characteristics. In this paper, we refer to these kinds of unorthodox filters as low/high-pass gain filters. Specifically, the momentum system behaves as a low-pass gain filter when $u_{k}>0,v_{k}=1$ and a high-pass gain filter when $u_{k}<0,v_{k}=1$. Additionally, if $u_{k}$ is close to $1$, the momentum system attenuates high-frequency gradient components while strongly amplifying low-frequency components; if $u_{k}$ is close to $-1$, the momentum system attenuates low-frequency gradient components while strongly amplifying high-frequency components. The visualization of the (dynamic) magnitude responses of unorthodox momentum systems is in Section 3.2 and Appendix B.2.
To demonstrate the momentum effects from the frequency perspective, in Fig. 1, we compare an original sinusoidal signal, a noisy version injected with Gaussian noise, and the signal after applying the momentum method (which is called momentum signal for short) in the time domain. The red curve represents the noisy signal, the black dashed curve corresponds to the original noise-free true signal, and the cyan curve shows the momentum signal. We can see that different selections of $u_{k}$ and $v_{k}$ significantly affect the amplifying or attenuating effects of the momentum system.
@@
Phase Response. We also have the phase response of the momentum system in the $k$-th stage,
While the phase response of the momentum only provides limited insights, understanding the behavior of the magnitude response across stages is essential for analyzing the time-variant characteristics of the momentum system. By plotting the dynamic magnitude response value $|H_{k}(\omega)|$ on the normalized angular frequency axis for each stage $k$, we can track how the frequency-dependent behavior of the multistage momentum system evolves over time. This provides valuable insights into the amplifying or attenuating characteristics of the momentum system. Further results on the comparisons of momentum systems with different dynamic magnitude responses are presented in the next section.
3 Dynamic Magnitude Response of the Momentum System
In this section, we present an empirical study to discover the influence of the momentum coefficients by comparing the test performance on momentum systems with different dynamic magnitude responses. We train VGG (Simonyan and Zisserman,, 2014) on the CIFAR-10 (Krizhevsky et al.,, 2009) dataset and ResNet50 (He et al.,, 2016) on the CIFAR-100 dataset using different momentum coefficients, while keeping all other hyperparameters unchanged. For each experiment, we report the mean and standard error (as subscripts) of test accuracy for 3 runs with random seeds from 0-2. The detailed experimental settings can be found in Appendix C. The experimental results in CIFAR-10 show high similarity to those in CIFAR-100. Thus, here, we mainly focus on the analysis based on CIFAR-100 and defer the experimental results of VGG16 on CIFAR-10 in Appendix B.3.
@@ 3.1 Orthodox Momentum Systems
Figure 2: (Up) Analysis of the (dynamic) magnitude responses in the early and late training stages for EMA-SGDM with low-pass momentum defined in Equation (10). The solid lines denote the magnitude responses in the early stages, and the dashed lines denote the magnitude responses in the late stages. (Down) The comparison between the gradient norms and momentum norms for EMA-SGDM with low-pass momentum. Left Column: increasing sequence. Middle Column: fixed sequence. Right Column: decreasing sequence.
@@
where $u_{t}\in[0,1)$ can be set as increasing, decreasing sequences, or fixed value. For time-variant momentum systems, different strategies of $u_{t}$ result in different time-variant filtering characteristics during training. According to Section 2.1, scaling the increasing and decreasing factors affects the changing rates of $u_{t}$. In the following, we demonstrate the dynamic magnitude responses, comparisons between gradient norms and momentum norms, and test accuracy results of orthodox momentum systems under different $u_{t}$ sequences  [^3] Note that selecting $\mu=100$ and $\nu=10^{4}$ lead to a long stage of super narrow-band filter. To avoid this problem, we select $\mu=10^{3},10^{4},10^{5}$ and $\nu=10^{2},10^{3},10^{4}$ in this paper..
Example 1: Low-Pass Momentum. We first explore the effect of increasing, fixed, and decreasing $u_{t}$ sequences in low-pass momentum. Fig. 2(a) - 2(c) show the corresponding dynamic magnitude responses over time. With increasing $u_{t}$, the system transits from an all-pass to a progressively narrower low-pass filter, gradually attenuating high-frequency components. Larger $\mu$ results in slower transitions. Decreasing $u_{t}$ shows a reverse behavior, with larger $\nu$ resulting in slower transitions. $u_{t}$ with a fixed value maintains a constant filter, with larger $u_{t}$ leading to more aggressive smoothing and noise reduction characteristics. The norm comparisons in Fig. 2(d) - 2(f) show that the momentum norms in low-pass momentum systems are always less than the corresponding gradient norms. Larger $u_{t},\nu$ and smaller $\mu$ lead to more reduced momentum norms, which validates the time-variant filtering characteristics of orthodox momentum systems.
@@
Example 2: High-Pass Momentum. High-pass momentum systems exhibit symmetric dynamic magnitude responses and similar norm comparisons, compared to their low-pass counterparts (see Fig. 6 in Appendix B.2). With increasing $u_{t}$, the system shifts from an all-pass to a narrow high-pass filter, progressively attenuating low-frequency components. Decreasing sequences act in reverse. Fixed sequences with larger $u_{t}$ lead to more aggressive attenuation of low-frequency components. The comparison of gradient norms and momentum norms can be found in Appendix B.2.
Test accuracy in Table 1 shows that dynamic high-pass systems with larger $\mu$ and smaller $\nu$ yield better top-1 accuracy performance. When selecting fixed values, momentum systems with larger $u_{t}$ perform more poorly. These results confirm that suppressing low-frequency gradient components is harmful. Moreover, high-pass systems generally outperform low-pass systems when applying decreasing strategies with the same $\nu$, suggesting that high-frequency components play a crucial role in the early training stages, which is also supported by the studies in Appendix B.4.
From Examples 1 and 2, we empirically verify that high-frequency gradient components are detrimental in late training stages, while their preservation in early stages leads to higher test accuracy, which matches the viewpoint that gradient noise has a generalization benefit early in training (Smith et al.,, 2020).

### v3

@@
Abstract
Momentum-based optimizers are widely adopted for training neural networks. However, the optimal selection of momentum coefficients remains elusive. This uncertainty impedes a clear understanding of the role of momentum in stochastic gradient methods. In this paper, we present a frequency domain analysis framework that interprets the momentum method as a time-variant filter for gradients, where adjustments to momentum coefficients modify the filter characteristics. Our experiments support this perspective and provide a deeper understanding of the mechanism involved. Moreover, our analysis reveals the following significant findings: high-frequency gradient components are undesired in the late stages of training; preserving the original gradient in the early stages, and gradually amplifying low-frequency gradient components during training both enhance performance. Based on these insights, we propose Frequency Stochastic Gradient Descent with Momentum (FSGDM), a heuristic optimizer that dynamically adjusts the momentum filtering characteristic with an empirically effective dynamic magnitude response. Experimental results demonstrate the superiority of FSGDM over conventional momentum optimizers.  [^1] Our implementation of FSGDM is available at https://github.com/yinleung/FSGDM.
[…]
@@ 1 Introduction
Momentum has achieved great success in deep learning applications when combined with Stochastic Gradient Descent (SGD) (Robbins & Monro, 1951). Among various momentum methods (Polyak, 1964; Nesterov, 1983; Van Scoy et al., 2017; Ma & Yarats, 2018; Kidambi et al., 2018), one of the most prevalent variants is the momentum method utilized within Stochastic Gradient Descent with Momentum (SGDM) (Sutskever et al., 2013; Paszke et al., 2019), which can be expressed as:
@@
where $g_{t}$ denotes the gradient at iteration $t$, ${m}_{t}$ is the momentum buffer, and $x_{t}$ represents the learnable parameters. The momentum coefficients $u_{t}$ and $v_{t}$ control the influence of the previous momentum and the current gradient, respectively, and $\alpha_{t}$ is the learning rate. For these time-variant momentum coefficients, a multistage setting has been commonly adopted in the machine learning community (Aybat et al., 2019; Kulunchakov & Mairal, 2019; Liu et al., 2020). Throughout this paper, we refer to this formulation, which decouples the two momentum coefficients, as Standard-SGDM. In contrast, another prevalent variant couples the two momentum coefficients using the Exponential Moving Average (EMA) method (Gardner Jr, 1985), leading to the formulation of EMA-SGDM:
@@
where $u_{t}\in[0,1)$ is the momentum coefficient. Notably, this coupled momentum formulation is a special case of the decoupled one, i.e., Standard-SGDM with $v_{t}=1-u_{t}$. Our experiments show performance gaps between these two formulations. Moreover, how the momentum coefficients change over time can significantly affect the test accuracy (see Section 3). The existence of these two distinct momentum formulations and their differing performances raise two primary questions in modern deep learning:
@@
For Question 1, some literature has investigated the convergence of the coupled method (Mai & Johansson, 2020; Li et al., 2022). Liu et al. (2020) argued that coupling the coefficients leads only to a constant scaling difference. Wang et al. (2024) further demonstrated that the mathematical equivalence between EMA-SGDM and Standard-SGDM can be achieved by adjusting the momentum coefficients and the learning rates in a coupled way. However, in practice, learning rate schedules are typically independent of momentum coefficient tuning during network training. On the other hand, popular frameworks like PyTorch (Paszke et al., 2019) adopt a decoupled momentum strategy by default. In our framework, we tackle the first question from the frequency domain perspective, revealing the relationship between the coupled and decoupled constructions.
@@
Regarding Question 2, prior research offered diverse opinions on how the momentum coefficients should vary over time. Some studies preferred fixed decoupled momentum coefficients (Yan et al., 2018; Liu et al., 2018; Yu et al., 2019), commonly selecting $u_{t}$ values as 0.9 and $v_{t}$ value as 1. Liu et al. (2020) highlighted the benefits of stagewise learning rate schedules in EMA-SGDM, noting that $u_{t}$ can either remain constant or increase along with the stagewise adjustments. Conversely, Smith (2018) demonstrated that decreasing the momentum coefficients while increasing the learning rate improves test performance. Moreover, Adaptive momentum methods (Kingma & Ba, 2014; Reddi et al., 2018; Luo et al., 2019; Chen et al., 2018) proved the convergence of decreasing coupled momentum coefficients in the context of online convex optimization. Nonetheless, a consensus regarding the optimal time-variant pattern of the momentum coefficients has yet to be reached.
@@
To answer these questions, one has to understand how the momentum method affects the training process. Goh (2017) analyzed the momentum method from the aspect of convergence and dynamics. Several prior studies (Cutkosky & Orabona, 2019; Ma & Yarats, 2018) speculated that averaging past stochastic gradients through momentum might reduce the variance of the noise in the parameter update, thus making the loss decrease faster. Polyak (1964); Rumelhart et al. (1986) argued that the EMA momentum can cancel out oscillations along high-curvature directions and add up contributions along low-curvature directions. From the signal processing perspective, the EMA method acts as a discrete low-pass filter for smoothing out high-frequency fluctuations while retaining the low-frequency baseband pattern of the signal (Gardner Jr, 1985). These points of view bring us a new insight into connecting the momentum update processes with the specific filters. In this aspect, the momentum methods with different coefficient selections can be interpreted in a unified frequency domain analysis framework, whereby Questions 1 and 2 are resolved.
@@
In this paper, we propose a novel frequency domain analysis framework to address the two questions and provide a deeper understanding of the role of momentum in stochastic optimization. To the best of our knowledge, this paper, for the first time, reveals the fundamental difference between Standard-SGDM and EMA-SGDM and uncovers the effects of the dynamic momentum coefficients clearly from the frequency domain perspective. This perspective not only explains the difference between various momentum methods but also provides practical guidelines for designing efficient optimizers. Accordingly, we introduce FSGDM, an optimizer that dynamically adjusts momentum filter characteristics during training. Experiments show that FSGDM outperforms traditional SGD-based momentum optimizers.
@@ 2 Frequency Domain Analysis Framework
This section introduces the background of Z-transform (Zadeh, 1950) in signal processing and then proposes a new frequency domain analysis framework for momentum methods.
@@ 2.1 Z-Transform and Quasi-Stationary Approximation
Frequency analysis is a crucial technique for understanding how systems react to varying frequency components of input signals. Specifically, for discrete-time linear time-invariant systems, Z-transform is leveraged to examine how systems attenuate or amplify signals at specific frequencies, especially in the study of system stability, pole-zero behavior, etc. (Oppenheim et al., 1996).
@@
However, one key issue is that the momentum system can be inherently time-variant, as its coefficients may change stagewise throughout the training process. This variability makes it difficult to apply traditional Z-transform analysis. To overcome this, inspired by the Zadeh (1961); Jury (1964), we approximate the system as time-invariant in each discrete interval stage. By holding the momentum coefficients constant over every interval, we construct a time-invariant quasi-stationary system (Hubner & Tran-Gia, 1991), enabling us to apply the Z-transform validly.
@@
where $\mu$ and $\nu$ are the increasing and decreasing factors  [^2] Note that different from the increasing sequence, the numerator of the decreasing sequence is $t+1$. This design avoids the zero gradients at the first training stage.. In Appendix C.1, we also examined the test set performance using other kinds of dynamic sequences. Under the above settings, for a given stage $k$ ($k=1,\cdots,N$), with $t\in[(k-1)\delta,k\delta-1]$, the momentum system becomes:
@@ 2.2 Frequency Domain Analysis of the Momentum Method
In this subsection, we introduce our frequency domain analysis framework and analyze the impacts of the momentum method on neural network training. We first apply Z-transform, denoted by $\mathcal{Z}$, to Equation 5:
@@
where $G(z)=\mathcal{Z}\{g_{t}\}$, $M(z)=\mathcal{Z}\{m_{t}\}$, and $z^{-1}M(z)=\mathcal{Z}\{m_{t-1}\}$. To obtain the frequency response of the momentum system during stage $k$, we evaluate the transfer function $H_{k}(z)$ on the unit circle (Oppenheim et al., 1996):
@@
Orthodox Momentum Systems are the ones whose amplitude of the magnitude response will not surpass $1$, like the EMA-SGDM (2). This kind of momentum system only shows attenuating characteristics. Specifically, the momentum system behaves as a low-pass filter when $u_{k}>0$ and a high-pass filter when $u_{k}<0$. Additionally, when $u_{k}$ gets close to 1, the momentum system will prefer to attenuate the gradient components with high frequencies. The visualization of the (dynamic) magnitude responses of orthodox momentum systems is in Section 3.1 and Appendix C.2.
For Unorthodox Momentum Systems where the amplitude of magnitude response will surpass $1$, such as selecting $u_{t}=0.9$ and $v_{t}=1$ in Standard-SGDM (1), the momentum system possesses both amplifying and attenuating characteristics. In this paper, we refer to these kinds of unorthodox filters as low/high-pass gain filters. Specifically, the momentum system behaves as a low-pass gain filter when $u_{k}>0,v_{k}=1$ and a high-pass gain filter when $u_{k}<0,v_{k}=1$. Additionally, if $u_{k}$ is close to $1$, the momentum system attenuates high-frequency gradient components while strongly amplifying low-frequency components; if $u_{k}$ is close to $-1$, the momentum system attenuates low-frequency gradient components while strongly amplifying high-frequency components. The visualization of the (dynamic) magnitude responses of unorthodox momentum systems is in Section 3.2 and Appendix C.2.
To demonstrate the momentum effects from the frequency perspective, in Figure 1, we compare an original sinusoidal signal, a noisy version injected with Gaussian noise, and the signal after applying the momentum method (which is called momentum signal for short) in the time domain. The red curve represents the noisy signal, the black dashed curve corresponds to the original noise-free true signal, and the cyan curve shows the momentum signal. We can see that different selections of $u_{k}$ and $v_{k}$ significantly affect the amplifying or attenuating effects of the momentum system.
@@
Similarly, we also have the phase response of the momentum system (see Appendix A). While the phase response of the momentum only provides limited insights, understanding the behavior of the magnitude response across stages is essential for analyzing the time-variant characteristics of the momentum system. By plotting the dynamic magnitude response value $|H_{k}(\omega)|$ on the normalized angular frequency axis for each stage $k$, we can track how the frequency-dependent behavior of the multistage momentum system evolves. This provides valuable insights into the amplifying or attenuating characteristics of the momentum system. Further results on the comparisons of momentum systems with different dynamic magnitude responses are presented in the next section.
3 Dynamic Magnitude Response of the Momentum Systems
In this section, we present an empirical study to discover the influence of the momentum coefficients by comparing the test performance on momentum systems with different dynamic magnitude responses. We train VGG (Simonyan & Zisserman, 2014) on the CIFAR-10 (Krizhevsky et al., 2009) dataset and ResNet50 (He et al., 2016) on the CIFAR-100 dataset using different momentum coefficients, while keeping all other hyperparameters unchanged. For each experiment, we report the mean and standard error (as subscripts) of test accuracy for 3 runs with random seeds from 0-2. The detailed experimental settings can be found in Appendix D. The experimental results in CIFAR-10 show high similarity to those in CIFAR-100. Thus, here, we mainly focus on the analysis based on CIFAR-100 and defer the experimental results of VGG16 on CIFAR-10 in Appendix C.3.
@@ 3.1 Orthodox Momentum Systems
Figure 2: (Up) Analysis of the (dynamic) magnitude responses in the early and late training stages for EMA-SGDM with low-pass momentum defined in Equation 9. The solid lines denote the magnitude responses in the early stages, and the dashed lines denote the magnitude responses in the late stages. (Down) The comparison between the gradient norms and momentum norms for EMA-SGDM with low-pass momentum. Left Column: increasing sequence. Middle Column: fixed sequence. Right Column: decreasing sequence.
@@
where $u_{t}\in[0,1)$ can be set as increasing, decreasing sequences, or fixed value. For time-variant momentum systems, different strategies of $u_{t}$ result in different time-variant filtering characteristics during training. According to Section 2.1, scaling the increasing and decreasing factors affects the changing rates of $u_{t}$. In the following, we demonstrate the dynamic magnitude responses, comparisons between gradient norms and momentum norms, and test accuracy results of orthodox momentum systems under different $u_{t}$ sequences  [^3] Note that selecting $\mu=100$ and $\nu=10^{4}$ lead to a long stage of the super narrow-band filter. To avoid this problem, we select $\mu=10^{3},10^{4},10^{5}$ and $\nu=10^{2},10^{3},10^{4}$ in this paper..
Example 1: Low-Pass Momentum. We first explore the effect of increasing, fixed, and decreasing $u_{t}$ sequences in low-pass momentum. Figure 2(a) - 2(c) show the corresponding dynamic magnitude responses over time. With increasing $u_{t}$, the system transits from an all-pass to a progressively narrower low-pass filter, gradually attenuating high-frequency components. Larger $\mu$ results in slower transitions. Decreasing $u_{t}$ shows a reverse behavior, with larger $\nu$ resulting in slower transitions. $u_{t}$ with a fixed value maintains a constant filter, with larger $u_{t}$ leading to more aggressive smoothing and noise reduction characteristics. The norm comparisons in Figure 2(d) - 2(f) show that the momentum norms in low-pass momentum systems are always less than corresponding gradient norms. Larger $u_{t},\nu$ and smaller $\mu$ lead to more reduced momentum norms, which validates the time-variant filtering characteristics of orthodox momentum systems.
@@
Example 2: High-Pass Momentum. High-pass momentum systems exhibit symmetric dynamic magnitude responses and similar norm comparisons, compared to their low-pass counterparts (see Figure 6 in Appendix C.2). With increasing $u_{t}$, the system shifts from an all-pass to a narrow high-pass filter, progressively attenuating low-frequency components. Decreasing sequences act in reverse. Fixed sequences with larger $u_{t}$ lead to more aggressive attenuation of low-frequency components. The comparison of gradient norms and momentum norms can be found in Appendix C.2.
Test accuracy in Table 1 shows that dynamic high-pass systems with larger $\mu$ and smaller $\nu$ yield better top-1 accuracy performance. When selecting fixed values, momentum systems with larger $u_{t}$ perform more poorly. These results confirm that suppressing low-frequency gradient components is harmful. Moreover, high-pass systems generally outperform low-pass systems when applying decreasing strategies with the same $\nu$, suggesting that high-frequency components play a crucial role in the early training stages, which is also supported by the studies in Appendix C.4.
From Examples 1 and 2, we empirically verify that high-frequency gradient components are detrimental in late training stages, while their preservation in early stages leads to higher test accuracy, which matches the viewpoint that gradient noise has a generalization benefit early in training (Smith et al., 2020).